I am thinking of 3 consecutive positive numbers. If I multiply the first with the third and then add the second, the result is 41. What are the numbers?

Answer by checkley71(8403)

let x, x+1 & x+2 be the 3 consecutive numbers.
x(x+2)+(x+1)=41
x^2+2x+x+1=41
x^2+3x+1-41=0
x^2+3x-40=0
(x+8)(x-5)=0
x-5=0
x=5 answer for the smaller number
5+1=6 for the next number
5+2=7for the largest number
proof
5*7+6=41
35+6=41
41=41